OpenStudy (anonymous):

Choose two axioms that allow 6 + (x + 5) to be written x + 11. commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

commutative - addition distributive

OpenStudy (anonymous):

associative and commutative addition

OpenStudy (anonymous):

distributive involves distributing to the whole problem best example is 4(x+2)=4x+8 the 4 is distributed to the x and the 2

OpenStudy (anonymous):

associative means that you can move the parenthesis around the problem and still get the same answer. like (4+2)+3=4+(2+3)

jhonyy9 (jhonyy9):

first and 6th is true